# GeoTexDensifier: Geometry-Texture-Aware Densification for High-Quality Photorealistic 3D Gaussian Splatting

### v1

@@ VII Experiments
Room (indoor) 1.54 0.919 31.414 0.219 0.62 0.901 29.908 0.251 2.57 0.922 31.579 0.208 3.93 0.928 31.576 0.187 2.73 0.927 31.791 0.196
Counter 1.21 0.909 28.993 0.2 0.56 0.889 27.899 0.235 2.57 0.915 29.186 0.183 3.78 0.913 28.728 0.171 1.74 0.916 29.216 0.179
Bonsai 1.26 0.942 31.991 0.203 0.76 0.925 31.026 0.225 2.08 0.947 32.554 0.191 3.76 0.947 31.957 0.174 1.83 0.954 32.952 0.161
Kitchen 1.82 0.926 31.036 0.127 0.73 0.908 29.207 0.148 3.19 0.931 31.702 0.119 3.67 0.933 31.802 0.114 2.08 0.93 31.547 0.12
Courtroom 2.99 0.818 23.621 0.217 1.47 0.797 23.152 0.265 5.18 0.818 23.719 0.215 5.24 0.792 21.73 0.219 4.78 0.829 24.287 0.199
Auditorium 0.69 0.881 24.083 0.247 0.25 0.867 23.765 0.292 1.21 0.883 24.351 0.244 4.8 0.839 21.506 0.274 1.64 0.892 25.24 0.234
Meeting Room 1.33 0.88 25.664 0.216 0.57 0.861 24.752 0.264 3.16 0.881 25.857 0.21 4.21 0.869 24.859 0.21 2.2 0.888 26.355 0.201
Church 2.24 0.829 22.946 0.247 1 0.811 22.355 0.279 3.61 0.833 23.317 0.243 4.86 0.818 21.89 0.241 4.08 0.845 23.678 0.217
Bicyle (outdoor) 6.1 0.765 25.205 0.21 2.8 0.75 24.891 0.237 9.06 0.778 25.265 0.181 6.02 0.798 25.542 0.158 6.03 0.782 25.388 0.177
Garden 5.88 0.866 27.34 0.107 2.91 0.858 27.117 0.121 8.78 0.871 27.492 0.098 5.82 0.877 27.517 0.091 5.36 0.87 27.561 0.104
Stump 4.92 0.773 26.621 0.215 1.93 0.753 25.502 0.247 6.66 0.786 26.842 0.187 5.37 0.804 27.114 0.169 4.93 0.786 26.874 0.195
Treehill 3.77 0.633 22.523 0.325 1.91 0.629 22.058 0.344 7.97 0.634 22.21 0.276 4.86 0.641 22.222 0.261 3.73 0.635 22.379 0.293
Flowers 3.63 0.606 21.578 0.336 2.09 0.578 21.065 0.371 7.48 0.635 21.546 0.262 4.87 0.643 21.528 0.254 6.18 0.621 21.251 0.26
Wufeng Tower 4.69 0.775 23.073 0.292 1.62 0.729 21.351 0.355 11.98 0.777 22.848 0.267 5.86 0.768 22.495 0.273 10.34 0.8 23.402 0.232
Office Park 2.09 0.684 22.96 0.42 0.53 0.643 22.588 0.477 1.82 0.668 22.991 0.448 7.16 0.527 15.413 0.557 7.02 0.754 24.309 0.327
We evaluate our GeoTexDensifier on eight indoor scenes and seven outdoor scenes, including nine cases from from Mip-NeRF 360 [6], four indoor cases “Church”, “Auditorium”, “Courtroom” and “Meeting Room” from Tanks and Temples [16], and two real outdoor scenes “Wufeng Tower” and “Office Park” captured by DJI Zenmuse P1 and Phantom 4 RTK respectively, each of which is composed of multi-view digital images, corresponding camera intrinsic and extrinsic parameters, and sparse SfM map points recovered by COLMAP [39]. Each experimental case typically contains some fully textured regions with appearance details, and some weakly textured surfaces. We implement GeoTexDensifier based on the source code of original 3DGS [1] and compare our implementation with original 3DGS and three other SOTA methods Mini-Splatting-D [14], GeoGaussian [15] and Pixel-GS [17]. All the results are trained with the source code of each method conducted on a single Ubuntu18.04 server with an Nvidia GeForce RTX4090 GPU with $24$GB memory, and $500$GB RAM, with the metrics SSIM, PSNR and LPIPS given in subsection VII-A to evaluate the corresponding rendering quality. We also exhibit ablation studies on texture-aware densification and geometry-aware splitting modules in subsection VII-B to show the usefulness of each strategy on the improvements of the reconstruction quality of 3DGS models.
@@ VII-A Qualitative and Quantitative Evaluation
Fig. 1 already gives an indoor scene “Church” for comparison to SOTA works. As the comparison results of other cases shown in Fig. 6 and Table I, the original 3DGS [1] often lacks sufficient splats in fully textured areas due to its simple densification strategy. Pixel-GS [17] and Mini-Splatting-D [14] address this issue to some extent, by counting the contribution of each Gaussian in two different strategies to find under-densified splats for further splitting. However, Pixel-GS still struggles to recover adequate details in the subtle textures such as the carpet in Fig. 1 and the chair in case “Meeting Room”, and seems to perform even worse in texture detail recovery than original 3DGS in self-captured case “Office Park” as shown in Fig. 6. This under-densification issue might still occur because Pixel-GS simply regards the covered pixel numbers as the weights for averaging pixel gradients, which turns out to be less adaptable to complex scenes than original strategy. On the other hand, Mini-Splatting-D employs a more explicit strategy to enforce splitting for reconstructing more texture details. Nevertheless, it suffers from over-splitting issue which further degrades reconstruction quality in areas with weak textures or insufficient training views. GeoGaussian introduces a planar regularization term to optimize the scene structure, but heavily relies on the quality of the input point cloud. Therefore, it specially uses PlanarSLAM [41] to obtain a more evenly distributed point cloud with high-quality even in weakly textured areas, which might not be suitable for general purpose. When using sparse SfM map points as input in our experiment for fair comparison, its performance is deteriorated significantly by failing to provide sufficient splats in texture-rich areas as well. In comparison, our GeoTexDensifier fully combines texture-aware densification and geometry-aware splitting strategies to produce a 3DGS model with cleaner, more reasonably distributed splats, which offers superior rendering quality compared to the other works as verified by the quantitative evaluation of metrics SSIM, PSNR and LPIPS given in Table I. Please refer to the supplementary material for more comparison details in the videos.
@@ VII-B Ablation Studies
VII-C VR Application
We can also import the 3DGS model generated by our GeoTexDensifier into an open-sourced “Gaussian Splatting VR Viewer Unity Native Plugin”  [^1] https://github.com/clarte53/GaussianSplattingVRViewerUnity to achieve photorealistic immersive visualization and real-time interactive navigation of the reconstructed scene on an Oculus Quest II standalone VR headset equipped with a single Ubuntu18.04 desktop PC with Intel Core i7-8700 CPU @ 3.2GHz, an Nvidia GeForce RTX4090 GPU, and $48$GB RAM, as illustrated in Fig. 8. Note that the example of the “Wufeng Tower” can be immersively navigated by 3DGS representation with highly restored geometry and texture details. Please refer to the supplementary video for the complete demo.
@@ VIII Conclusion and Future Work
Like other SOTA methods, our approach introduces more splats to reconstruct finer details in textured areas, which results in higher GPU memory consumption than original 3DGS when handling large-scale scenes. We plan to extend our method to support urban-scale reconstruction and navigation by integrating techniques such as chunk-based reconstruction, LOD organization and dynamically scheduled rendering. Additionally, the potential possibility of combining our approach with generative diffusion models will be explored as a future work to further enhance reconstruction completion and quality in insufficiently observed or invisible regions.
@@ Acknowledgments
The authors wish to thank Zhuang Zhang, Chongshan Sheng, Fei Jiao, Xia Sun, Jinyi Liao, Renzhi Wang and Yuqing Xie for their kind helps in the development and the experiments of the proposed GeoTexDensifier system. This work was partially supported by NSF of China (No. 62425209).

### v2

@@ VII Experiments
Bicycle 5.49 0.771 25.262 0.199 6.28 0.777 25.372 0.19 5.73 0.78 25.367 0.186 5.81 0.78 25.407 0.187 6.03 0.782 25.388 0.177
Garden 5.33 0.868 27.433 0.105 5.91 0.869 27.536 0.105 5.36 0.87 27.541 0.103 5.37 0.87 27.555 0.103 5.36 0.87 27.561 0.104
Stump 4.82 0.776 26.672 0.209 4.97 0.783 26.775 0.196 5.02 0.786 26.866 0.195 4.67 0.786 26.869 0.196 4.93 0.786 26.874 0.195
Treehill 3.43 0.663 22.367 0.324 4.14 0.631 22.282 0.29 3.63 0.633 22.315 0.293 3.8 0.635 22.422 0.292 3.73 0.635 22.379 0.293
Flower 3.51 0.607 21.478 0.334 7.74 0.613 21.132 0.261 6.29 0.621 21.23 0.259 6.38 0.621 21.21 0.258 6.18 0.621 21.251 0.26
Room 1.43 0.919 31.455 0.218 3.82 0.926 31.493 0.193 2.52 0.927 31.713 0.196 2.67 0.927 31.755 0.194 2.55 0.927 31.93 0.195
Counter 1.06 0.907 28.88 0.201 2.29 0.915 29.183 0.177 1.7 0.915 29.178 0.18 1.74 0.915 29.191 0.18 1.74 0.916 29.216 0.179
Bonsai 1.1 0.948 32.477 0.207 2.15 0.954 32.947 0.158 1.81 0.954 32.674 0.161 1.84 0.954 32.95 0.161 1.83 0.954 32.952 0.161
Kitchen 1.39 0.926 31.138 0.127 2.68 0.931 31.641 0.118 2.12 0.931 31.463 0.119 2.1 0.931 31.331 0.119 2.09 0.931 31.698 0.12
We evaluate our GeoTexDensifier on nine cases from Mip-NeRF 360 [6], twenty one cases from Tanks and Temples [16], and two real outdoor scenes “Wufeng Tower” and “Office Park” captured by DJI Zenmuse P1 and Phantom 4 RTK respectively, each of which is composed of multi-view digital images, corresponding camera intrinsic and extrinsic parameters, and sparse SfM map points recovered by COLMAP [39]. Each experimental case typically contains some fully textured regions with appearance details, and some weakly textured surfaces. We implement GeoTexDensifier based on the source code of original 3DGS [1] and compare our implementation with original 3DGS, original 3DGS with depth prior from Depth Anything v2 (DAv2) [41], and three other SOTA methods Mini-Splatting-D [14], GeoGaussian [15] and Pixel-GS [17]. All the results are trained with the source code of each method conducted on a single Ubuntu18.04 server with an Nvidia GeForce RTX4090 GPU with 2424GB memory, and 500500GB RAM, with the metrics SSIM, PSNR and LPIPS given in subsection VII-A to evaluate the corresponding rendering quality. We also exhibit ablation studies on key modules such as texture-aware densification, geometry-aware splitting with VDRC and normal guidance in subsection VII-B to show the usefulness of each strategy on the improvements of the reconstruction quality of 3DGS models.
@@ VII-A Qualitative and Quantitative Evaluation
Fig. 1 already gives an indoor scene “Church” for comparison to SOTA works. As the comparison results of other cases shown in Fig. 6 and Table II, the original 3DGS [1] often lacks sufficient splats in fully textured areas due to its simple densification strategy. Pixel-GS [17] and Mini-Splatting-D [14] address this issue to some extent, by counting the contribution of each Gaussian in two different strategies to find under-densified splats for further splitting. However, Pixel-GS still struggles to recover adequate details in the subtle textures such as the carpet in Fig. 1 and the chair in case “Meeting Room”, and seems to perform even worse in texture detail recovery than original 3DGS in self-captured case “Office Park” as shown in Fig. 6. This under-densification issue might still occur because Pixel-GS simply regards the covered pixel numbers as the weights for averaging pixel gradients, which turns out to be less adaptable to complex scenes than original strategy. On the other hand, Mini-Splatting-D employs a more explicit strategy to enforce splitting for reconstructing more texture details. Nevertheless, it suffers from over-splitting issue which further degrades reconstruction quality in areas with weak textures or insufficient training views. GeoGaussian introduces a planar regularization term to optimize the scene structure, but heavily relies on the quality of the input point cloud. Therefore, it specially uses PlanarSLAM [42] to obtain a more evenly distributed point cloud with high-quality even in weakly textured areas, which might not be suitable for general purpose. When using sparse SfM map points as input in our experiment for fair comparison, its performance is deteriorated significantly by failing to provide sufficient splats in texture-rich areas as well. Original 3DGS with depth prior is more similar to our geometry-aware splitting strategy, and outperforms the original one without depth prior. However, due to the lack of true scale for DAv2 [41], the source code implementation of original 3DGS aligns the DAv2 depth to the sparse SfM points to obtain a scale and offset for each frame as depth prior, whose accuracy might be affected by some unreliable SfM points or possible scale inconsistency across multiple frames to impact the final rendering quality. In comparison, our GeoTexDensifier fully combines texture-aware densification and geometry-aware splitting strategies to produce a 3DGS model consisting of cleaner, more reasonably and more accurately distributed splats without rely on truly scaled depth prior, which offers superior rendering quality compared to the other works as verified by the quantitative evaluation of metrics SSIM, PSNR and LPIPS given in Table II. Please refer to the supplementary material for more comparison details in the videos.
@@ VII-B Ablation Studies
Table III further provides more quantitative comparisons of ablation strategies including only geometry-aware splitting (Geo.), only texture-aware densification (Tex.), texture-aware densification with VDRC (depth prior) (Tex. + VDRC), texture-aware densification with normal guidance (Tex. + Normal), and GeoTexDensifier (full strategies with Tex. + Geo.), on Mip-NeRF 360 dataset [6]. We can see from the quatitative evaluation metrics with SSIM, PSNR and LPIPS in the ablation experiments that texture-aware densification strategy makes greater improvement in 3DGS quality than using only depth prior or normal prior due to better distributed splats, while texture-aware densification combined with either depth prior or normal prior further improves the reconstruction quality with more accurate optimization of splat positions in textureless areas. Finally, full strategies combining both depth prior and normal prior with texture-aware densification can make the best performance in the final rendering quality.
VII-C Limitations
Our geometry-aware splitting strategy effectively resolves weakly-textured issues by filtering erroneous Gaussian splats via VDRC checking. However, this strategy might introduce a limitation in object-centric scenarios captured with distant backgrounds such as the two outdoor cases of “Horse” and “Truck” from Tanks and Temples dataset [16] shown in Fig. 8. Initial SfM points for such distant regions are usually sparse or missing. When Gaussian splats in foreground regions undergo splitting, the large relative depth ratio changes on foreground/background boundary regions typically fail to pass the VDRC checking, which consequently prevent these splats to be split to the distant regions, leading to “under-reconstruction” regions with missing Gaussians. These incomplete regions possibly cause blurs or ghosting artifacts in the distant backgrounds from certain viewpoints, as exemplified in the magnified regions of the representative views of the two cases in Fig. 8.
VII-D VR Application
We can also import the 3DGS model generated by our GeoTexDensifier into an open-sourced “Gaussian Splatting VR Viewer Unity Native Plugin”  [^1] https://github.com/clarte53/GaussianSplattingVRViewerUnity to achieve photorealistic immersive visualization and real-time interactive navigation of the reconstructed scene on an Oculus Quest II standalone VR headset equipped with a single Ubuntu18.04 desktop PC with Intel Core i7-8700 CPU @ 3.2GHz, an Nvidia GeForce RTX4090 GPU, and 4848GB RAM, as illustrated in Fig. 9. Note that the example of the “Wufeng Tower” can be immersively navigated by 3DGS representation with highly restored geometry and texture details. Please refer to the supplementary video for the complete demo.
@@ VIII Conclusion and Future Work
Our limitation in object-centric scenes with distant backgrounds discussed in section VII-C can be better solved by combining a more reliable dense point cloud initialization instead of sparse SfM points to further avoid “under-reconstruction” situation. Besides, like other SOTA methods, our approach introduces more splats to reconstruct finer details in textured areas, which results in higher GPU memory consumption than original 3DGS when handling large-scale scenes. Our VDRC checking and normal guidance operations also increase the training time complexity. We plan to extend our method to support urban-scale reconstruction and navigation in a more space/time-efficient way by integrating techniques such as chunk-based reconstruction, LOD organization and dynamically scheduled rendering. Additionally, the potential possibility of combining our approach with generative diffusion models will be explored as a future work to further enhance reconstruction completion and quality in insufficiently observed or invisible regions.
@@ Acknowledgments
The authors wish to thank Zhuang Zhang, Chongshan Sheng, Fei Jiao, Xia Sun, Jinyi Liao, Renzhi Wang, Yuqing Xie, Bingze Li and Hongliang Sun for their kind helps in the development and the experiments of the proposed GeoTexDensifier system. This work was partially supported by Key R&D Program of Zhejiang Province (No. 2023C01039), and NSF of China (No. 62425209).